Solve the differential equation 2y'''+ y''- 8y' - 4y = 7 - 2e-x - cos2 (2x).


1
Expert's answer
2021-08-11T19:20:45-0400

The homogeneous equation


"2y'''+y''-8y'-4y=0"

The characteristic equation


"2r^3+r^2-8r-4=0"

"r^2(2r+1)-4(2r+1)=0"

"(2r+1)(r+2)(r-2)=0"

"r_1=-\dfrac{1}{2}, r_2=-2, r_3=2"

The general solution of the homogeneous differential equation is


"y_h=C_1e^{-x\/2}+C_2e^{-2x}+C_3e^{2x}"


"7-2e^{-x}-\cos^2(2x)=7-2e^{-x}-\dfrac{1}{2}-\dfrac{1}{2}\cos(4x)"

"=\dfrac{13}{2}-2e^{-x}-\dfrac{1}{2}\cos(4x)"

Find the particular solution of the nonhomogeneous differential equation


"y_p=A+Be^{-x}+C\cos(4x)+D\sin(4x)"

"y_p'=-Be^{-x}-4C\sin(4x)+4D\cos(4x)"

"y_p''=Be^{-x}-16C\cos(4x)-16D\sin(4x)"

"y_p'''=-Be^{-x}+64C\sin(4x)-64D\cos(4x)"

Substitute


"-2Be^{-x}+128C\sin(4x)-128D\cos(4x)"

"+Be^{-x}-16C\cos(4x)-16D\sin(4x)"

"+8Be^{-x}+32C\sin(4x)-32D\cos(4x)"

"-4A-4Be^{-x}-4C\cos(4x)-4D\sin(4x)"

"=\dfrac{13}{2}-2e^{-x}-\dfrac{1}{2}\cos(4x)"

"A=-\dfrac{13}{8}"

"B=-\dfrac{2}{3}"

"D=8C"

"C=\dfrac{1}{2600}"

"D=\dfrac{1}{325}"

Then


"y_p=-\dfrac{13}{8}-\dfrac{2}{3}e^{-x}+\dfrac{1}{2600}\cos(4x)+\dfrac{1}{325}\sin(4x)"

The general solution of the nonhomogeneous differential equation is


"y=C_1e^{-x\/2}+C_2e^{-2x}+C_3e^{2x}"

"-\dfrac{13}{8}-\dfrac{2}{3}e^{-x}+\dfrac{1}{2600}\cos(4x)+\dfrac{1}{325}\sin(4x)"